OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x - 2) <= 2? A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left. 3 >= x x <= 3

OpenStudy (anonymous):

2(x-2) <= 2 2x - 4 <= 2 +4 +4 2x <= 6 2 2 x = 3 x >= 3

OpenStudy (anonymous):

A number line with a closed circle on 3 and shading to the right.

OpenStudy (anonymous):

that is correct ^^^... that is NOT a way to represent the solution

OpenStudy (anonymous):

ok thxs